# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ 5.4 Dataset Popularity
As outlined in this paper, the Network Intrusion Detection domain has long been plagued by the lack of reliable datasets. For this reason, many researchers have relied on DARPA-based datasets – KDD'99 and NSL-KDD, for over two decades. While many recent studies [177, 178, 179, 44] still consider them the most popular, such studies tend to analyze very extended periods, sometimes as far back as 2008 [44]. Therefore, despite KDD'99 and NSL-KDD remaining undoubtedly popular, we argue that analyzing such long time spans fails to capture contemporary domain trends properly.
For this reason, we conducted our own data popularity analysis of NID papers published between 2020 and 2023 at top-tier security conferences (details in A.2). In total, we identified 45 relevant papers and summarized their utilized datasets in Figure 4. Most papers evaluate using more than one dataset (often three or four), causing the total sum to be higher than 45. We summarize key findings below, with a complete list of papers and their analysis on our GitHub.
@@
Custom data to get the job done. This survey shows that while many datasets exist, most current research prefers to collect its own data (category Custom). However, not sharing them often renders the associated research non-replicable.
@@
Not a clear winner. Even if researchers decide to utilize public datasets, our results indicate their relatively evenly distributed usage. Although CIC-IDS2017 came out as the most popular, its utilization only in 29% of papers brings it far from being a “standardized” benchmark in the domain. In addition, most of its users rely on the original version with known errors, potentially undermining their research outcomes and suggesting their unawareness of the dataset’s flaws and fixes.
The other datasets exceeding five uses include CTU-13, Kitsune, and UNSW-NB15. Datasets listed in this survey (Table LABEL:tab:data_survey) but utilized only once were aggregated under the Other (listed) category, whereas Other (unlisted) groups datasets not present in the summary table. This category mainly refers to reused datasets from other papers that produced data as a research side-effect. However, it also involves datasets like MACCDC [181], published as a part of the NETRESEC data list, or CIC DoS 2017 [182], unavailable at the time of writing this paper. Refer to our GitHub for more details.
@@ 5.5 Trends in NID Dataset Research
This subsection analyzes trends in datasets and their properties based on Table LABEL:tab:data_survey. We primarily focus on property changes over time, as well as differences between new datasets (published since 2020) and older ones. In the analysis, we consider the data publication date rather than its collection year.
@@ 5.5.1 Increase in Number While Narrowing the Scope
Along with publishing more data, the datasets are also becoming more specialized. Since 2016, special-purpose datasets have dominated the field (83% of published datasets), a significant shift from the previous balance with general-purpose datasets. Concerning environment-specific datasets, IoT is the most popular environment, present in 15% (20% with IIoT and industrial networks) of all surveyed datasets. From the attacks’ perspective, most prevalent intrusion-specific datasets focus on botnet traffic (10%). Since 2020, four datasets focused on Advanced Persistent Threats (APT) have been released, emphasizing the growing significance of these threats.
@@ 5.5.2 Traffic Generation Strategies
Providing realistic normal and attack traffic should be a main focus of any dataset. As discussed throughout this paper, the network traffic can be obtained via real capture, emulation, and synthetic crafting – each with distinct advantages and drawbacks. This part elaborates on these techniques and explores their applications in existing datasets.
Normal traffic: While capturing normal traffic from a production network offers the most realism, privacy concerns and anonymization needs are often too discouraging. For this reason, only 17% of datasets provide purely real normal traffic, while an additional 8% mix it with emulation.
Aiming to overcome issues with real traffic, many datasets utilize emulation: 70% rely solely on it, while 77.5% include it alongside other traffic types. This trend is even stronger for new datasets – 84% include emulation since 2020. While physical emulation is currently the most common (42% of emulated datasets), virtual emulation (30%) and their combination (28%) are also used. Notably, virtual emulation is gaining more prominence (60% of emulated datasets since 2020) due to improved computing and virtualization technologies.
Besides the emulation environment (physical or virtual), normal traffic is also characterized by its generation method. Regarding this matter, we observed four common patterns:
@@
    Emulation via traffic generators: Traffic generators and performance testers are a straightforward way to simulate background activity. However, they typically produce overly simplistic (i.e., monotonous) traffic that fails to reflect real network complexity. This fact makes it easier for classifiers to separate such traffic from anomalies and malicious patterns, leading to an overestimation of their real effectiveness. Examples employing traffic generators include FLNET2023 [156], UNR-IDD [76], and GTCS [120], using hping3, IPerf, and Ostinato tools respectively. Further information and a list of 92 traffic generators are provided by Adeleke et al. [183].
@@
    Incorporating humans: The most realistic – yet most expensive – way of traffic emulation involves humans performing regular tasks on computers or interacting with the network (e.g., using IoT devices). Consequently, this approach also requires to specify and follow pre-defined user profiles to ensure traffic diversity. Datasets with emulated user behavior include CUPID [56] or CICIoT2023 [153].
@@
Although not commonly present in public datasets, synthetic traffic generation is another option to produce normal traffic without privacy concerns. In the past, synthetic traffic (narrowed to a problem of network traffic matrices) was generated using simple statistical methods [184]. However, the rise of generative artificial intelligence (GenAI) has improved its sophistication significantly.
Ring et al. [185] were the first to use Generative Adversarial Networks (GANs) to generate network flows synthetically. The method used a novel technique, IP2Vec [186], to learn meaningful vector representations of IP addresses by transforming them into a continuous feature space. Manocchio et al. [187] introduced FlowGAN, a model based on Manifold Guided GAN architecture to generate NetFlow v9 data with more realistic data distributions. While GANs are popular, the authors of STAN [188] propose using an autoregressive neural network with convolutional layers to capture realistic spatio-temporal network traffic characteristics.
Generated traffic, whether emulated or synthetic, should aim for a high degree of realism. Synthetic generation methods are often evaluated by statistical tests to compare their generated traffic distribution with other data, typically from the distribution the synthetic generator was trained on. For instance, Schoen et al. [52] demonstrated a strong validation approach using eight metrics, statistical tests, and domain knowledge to assess the realism, diversity, novelty, and compliance of the synthetically generated network flow data. As presented by Cha et al. [189], various other distance metrics for comparing probability density functions can also be applied.
@@
Similar to normal traffic, emulation was the most popular way for generating malicious traffic, used in 84% of all datasets (79% for sole emulation without other traffic types). This trend has strengthened in recent datasets, reaching 90% (86% for sole emulation). Physical emulation remains the most popular, with virtual emulation gaining in popularity. The most frequent attacks were DoS (74% of all datasets), Reconnaissance (67%), and Other (57%), rising to 65% for datasets published since 2020 due to the increase in special-purpose datasets.
In general, emulating malicious traffic is less complex than normal traffic, so rigorous statistical tests are usually omitted. Instead, physical and virtual attacks are simulated using real-world tools, such as in the Kali Linux suite [54]. Nevertheless, as the realism of these attacks is also crucial, emphasis should be placed on creating multiple attack scenarios and varying the attack parameters, as in the AIT Log Dataset [43].
@@
Synthetic attack traffic is rare in public datasets, yet its generation and utilization are well-studied. In general, it can be used in two ways: a) oversampling or b) injection. While all synthetic traffic must comply with network functionality (align its timing characteristics with other traffic and generate valid data – e.g., a network flow cannot contain zero packets), attacks must also account for interactions with other traffic. For instance, a successful DDoS attack would influence the timing characteristics of all simultaneous network communication or introduce packet drops. For this reason, ensuring data realism with synthetic network attacks requires a delicate approach.
@@
Unlike oversampling, which enlarges the sample size of existing classes, injecting synthetic attacks creates new classes not previously present in the data. This enables the simulation of attack scenarios not actually performed on the network, offering a cost-efficient way of benchmarking NIDSs, as crafting synthetic samples is considered cheaper than emulation. The Intrusion Detection Dataset Toolkit (ID2T) [196, 197] allows the injection of synthetic attacks into raw PCAP files consisting of background traffic while accounting for potential attack effects like packet drops or delays. Similar tools, such as FLAME (2008) [198] and MACE (2004) [199], were also proposed but are no longer maintained, thus not offering newer attacks.
@@ 5.5.3 Popular Formats and Feature Sets
As no standardized feature set for flow-based data in the domain exists, dataset authors typically engineer their own features. This causes incompatibilities and degrades generalizability. Nevertheless, CICFlowMeter [39], introduced in CIC-IDS2017, challenges this trend, as 31% of surveyed flow-based datasets used it, making it the most popular tool for feature extraction. It is also adopted by many non-CIC datasets such as GTCS [120], TII-SSRC-23 [154], and FLNET2023 [156], and has also been used to transform other existing datasets like ToN_IoT and Bot-IoT into its features [40].
Although a step towards feature standardization is positive in terms of generalizability and transferability, locking onto a single feature set can also bring downsides. Komisarek et al. [152] criticized CICFlowMeter features as unsuitable for real-time intrusion detection. Sarhan et al. [40] investigated the standardization of NetFlow v9 features for NID and stated that they outperform CIC-based features in attack detection.
@@ 5.5.4 Other Observations
From the temporal perspective, 33% of datasets are captured in a single continuous run, 18% are periodic, and 37% are discontinuous – gaps within the data exist – with the latter increasing to 47% since 2020. Similarly to simulation in small networks, we argue that discontinuous captures are easier to perform, allowing flexible adjustments and less strict scheduling. Therefore, they pose a convenient choice for data creators.
@@ 6 Recommendations: Data Selection, Creation, & Usage
After reviewing existing datasets, this section outlines best practices on data-related NIDS topics: selecting the most appropriate dataset for one’s needs (Section 6.1), creating custom datasets (Section 6.2), and proper data handling to mitigate biases during experimentation (Section 6.3).
@@ 6.1 Choosing a Dataset
Using a public dataset is the most straightforward and time-efficient approach for NIDS experimentation. As dozens of such datasets are available, the following paragraphs guide the selection of the most suitable one for specific use cases. However, we refrain from suggesting concrete datasets, as this survey does not aim to assess data quality but rather help others to make informed decisions.
@@
Quality estimation: While we do not assess dataset quality in this paper, we strongly recommend performing it as a part of the data selection process. Despite the previous research on the topic [203, 103], no standardized formal metric for NID data quality has been established within the domain [204]. As “quality” itself is vague, and user needs may vary, formalizing such a general metric is challenging. Therefore, we discuss data quality in a more practical sense and attempt to gauge it by analyzing a) the realism of the data based on the rigorousness of its creation process and b) the contents of the data itself.
A quality dataset should be realistic and error-free. As documentation (metadata) is one of the few ways to learn about how data was created, its analysis remains a main resource for estimating data realism. Its key details include data purpose, creation process, and related limitations (more desired details in Section 6.2.4). Given this information, datasets with robust traffic generation and realistic environments are preferred to more simplistic datasets. However, our experience shows that analyzing documentation alone is insufficient. Even with decent documentation, numerous datasets were found to contain discrepancies like incorrect entries or timestamps. For this reason, thorough data analysis (e.g., N/A values, incorrect timestamps, packet errors, labeling issues) is essential in addition to documentation analysis to gauge data correctness and realism.
@@ 6.2 Creating a Dataset
In essence, creating a dataset involves 10 steps, from defining objectives and designing scenarios to documenting and sharing the data, as depicted in Figure 6. In the following paragraphs, we elaborate on desirable properties and recommendations in individual data creation steps to mitigate common dataset limitations listed in Section 3.2.
Figure 6: The process of creating a NID dataset. The 10 steps are further grouped into four phases – preparation, data collection, post-processing, and publication. Ideally, the process is linear, and the data is collected and processed in one run. However, in practice, the process is often iterative, such as re-designing and adding new scenarios or collecting and post-processing the data in multiple turns.

### v2

@@ 5.4 Dataset Popularity
As outlined in this paper, the Network Intrusion Detection domain has long been plagued by the lack of quality datasets. For this reason, many researchers have relied on DARPA-based datasets – KDD'99 and NSL-KDD, for over two decades. While many recent studies [179, 180, 181, 46] still consider them the most popular, such studies tend to analyze extended time periods, sometimes dating back to 2008 [46]. While KDD'99 and NSL-KDD remain extensively used, we argue that analyzing across such long time spans fails to reflect contemporary trends in the field properly.
For this reason, we conducted our own data popularity analysis of NID papers published between 2020 and 2023 at top-tier security conferences (details in A.2). In total, we identified 45 relevant papers and summarized their utilized datasets in Figure 4. Most papers evaluate using more than one dataset (often three or four), causing the total sum to be higher than 45. We discuss key findings below, with a complete list of papers and their analysis on our GitHub.
@@
Custom data to get the job done: This survey shows that while many datasets exist, most current research prefers to collect its own data (category Custom). However, the lack of data sharing often renders the associated research non-replicable.
@@
Not a clear winner: Even if researchers decide to utilize public datasets, our findings indicate that their usage is relatively evenly distributed. While CIC-IDS2017 emerged as the most popular, its utilization only in 29% of papers brings it far from being a “standardized” benchmark in the domain. In addition, most of its users rely on the original version with known errors, potentially undermining their research outcomes and indicating their unawareness of the dataset’s flaws and fixes.
The other datasets exceeding five uses include CTU-13, Kitsune, and UNSW-NB15. Datasets listed in this survey (Table LABEL:tab:data_survey) but utilized only once were aggregated under the Other (listed) category, whereas Other (unlisted) groups datasets not included in the summary table. This category mainly refers to reused datasets from other papers that produced data as a research side-effect. However, it also includes datasets like MACCDC [183], published as a part of the NETRESEC data list, and CIC DoS 2017 [184], unavailable at the time of writing. For more details, refer to our GitHub.
@@ 5.5 Trends in NID Dataset Research
This subsection analyzes trends in datasets and their properties based on Table LABEL:tab:data_survey. We primarily focus on property changes over time, as well as differences between new datasets (published since 2020) and older ones. For this analysis, we consider the data publication date rather than its collection year.
@@ 5.5.1 Increase in Number While Narrowing the Scope
Along with publishing more data, the datasets are also becoming more specialized. Since 2016, special-purpose datasets have dominated the field (83% of published datasets), a significant shift from the previous balance with general-purpose ones. Among environment-specific datasets, IoT is the most popular environment, present in 15% of all datasets (20% with IIoT and industrial networks). From an attack perspective, most prevalent intrusion-specific datasets focus on botnet traffic (10%) and DoS/DDoS (9%). Additionally, since 2020, four Advanced Persistent Threat (APT) datasets have been released, emphasizing the growing significance of these threats.
@@ 5.5.2 Traffic Generation Strategies
Providing realistic normal and attack traffic should be a primary objective of every dataset. As discussed throughout this paper, the network traffic can be acquired through real capture, emulation, and synthetic generation – each with distinct advantages and drawbacks. This part elaborates on these techniques and explores their applications in existing datasets.
Normal traffic: While capturing normal traffic from a production network offers the most realism, the associated privacy concerns and anonymization needs can be highly discouraging. For this reason, only 17% of datasets provide purely real normal traffic, while an additional 8% combine it with emulation.
Aiming to overcome the issues with real traffic, many datasets utilize emulation: 71% rely solely on it, while additional 8% include it alongside other traffic types. This trend is even stronger for new datasets – 84% include emulation since 2020. While physical emulation is the most common (41% of emulated datasets), virtual emulation (32%) and their combination (27%), are also used. Notably, virtual emulation is gaining more prominence (60% of emulated datasets since 2020) due to improved computing and virtualization technologies.
In addition to the emulation environment (physical or virtual), normal traffic is also characterized by its generation method. In this regard, we observed four common patterns:
@@
    Emulation via traffic generators: Traffic generators and performance testers are a straightforward way to simulate background activity. However, they typically produce overly simplistic (e.g., monotonous) traffic that fails to reflect real network complexity. This fact makes it easier for detectors to separate such traffic from anomalies and malicious patterns, leading to an overestimation of their real capabilities. Examples employing traffic generators include FLNET2023 [157], UNR-IDD [77], and GTCS [121], using hping3, IPerf, and Ostinato tools respectively. Further information and a list of 92 traffic generators are provided by Adeleke et al. [185].
@@
    Incorporating humans: The most realistic – yet most expensive – way of traffic emulation involves humans performing regular tasks on computers or interacting with the network (e.g., using IoT devices). Consequently, this approach also requires to specify and follow pre-defined user profiles to ensure traffic diversity. Datasets with this emulated behavior include CUPID [60] or CICIoT2023 [154].
@@
Although not commonly present in public datasets, synthetic traffic generation is another option to produce network traffic without privacy concerns. In the past, synthetic traffic, often narrowed to a network traffic matrix modeling, was generated using simple statistical methods [186]. However, recent advancements in generative artificial intelligence (GenAI) have improved its sophistication significantly.
Ring et al. [187] were the first to use Generative Adversarial Networks (GANs) to generate network flows synthetically. Their method used a novel technique, IP2Vec [188], to learn meaningful vector representations of IP addresses by transforming them into a continuous feature space. Manocchio et al. [189] introduced FlowGAN, a model based on Manifold Guided GAN architecture to generate NetFlow v9 data with more realistic data distributions. In addition to GANs, the authors of STAN [190] proposed an autoregressive neural network with convolutional layers to capture realistic spatio-temporal network traffic characteristics.
Generated traffic, whether emulated or synthetic, should aim for a high degree of realism. Synthetic generation methods are often evaluated by statistical tests that compare the distribution of generated traffic with other data, typically from the distribution the synthetic generator was trained on. For instance, Schoen et al. [56] demonstrated a robust validation approach using eight metrics, statistical tests, and domain knowledge to assess the realism, diversity, novelty, and compliance of the synthetically generated network flow data. As presented by Cha et al. [191], various other distance metrics for comparing probability density functions can also be applied.
@@
Similar to normal traffic, emulation was the most popular way for generating malicious traffic, used in 85% of all datasets (80% for sole emulation without other traffic types). This trend has strengthened in recent datasets, reaching 90% (86% for sole emulation). Physical emulation remains the most popular, with virtual emulation gaining in popularity. The most frequent attacks were DoS (74% of all datasets), Reconnaissance (67%), and Other (57%), rising to 65% for datasets published since 2020 due to the increase in special-purpose datasets.
In general, emulating malicious traffic is less complex than normal traffic, so rigorous statistical tests are often omitted. Instead, both physical and virtual attacks are typically simulated using real-world tools, such as in the Kali Linux suite [58]. However, since the realism of these attacks remains crucial, efforts should focus on designing diverse attack scenarios and varying attack parameters, as in the AIT Log Dataset [45].
@@
Synthetic attack traffic is uncommon in public datasets, yet its generation and utilization are well-studied. In general, it can be used in two ways: a) oversampling or b) injection. While all synthetic traffic must comply with network functionality (align its timing characteristics with other traffic and generate valid data – e.g., a network flow cannot contain zero packets), attacks should also account for interactions with other network activity. For instance, a successful DDoS attack would affect the timing characteristics of all simultaneous communication or introduce packet drops. For this reason, ensuring data realism with synthetic network attacks requires a delicate approach.
@@
Unlike oversampling, which enlarges the sample size of existing classes, injecting synthetic attacks creates new classes not previously present in the data. This enables the simulation of attack scenarios that were never executed on the network, offering a cost-efficient way of benchmarking NIDSs, as crafting synthetic samples is typically cheaper than emulation. The Intrusion Detection Dataset Toolkit (ID2T) [198, 199] allows the injection of synthetic attacks into raw PCAP files consisting of background traffic while accounting for potential attack effects like packet drops or delays. Similar tools, including FLAME (2008) [200] and MACE (2004) [201], were also developed but are no longer maintained, thus not supporting newer attacks.
@@ 5.5.3 Popular Formats and Feature Sets
As no standardized feature set for flow-based data in the domain exists, dataset authors typically engineer their own features. This causes incompatibilities and degrades generalizability. Nevertheless, CICFlowMeter [41], introduced in CIC-IDS2017, challenges this trend, as 31% of surveyed flow-based datasets used it, making it the most popular tool for feature extraction. It is also adopted by many non-CIC datasets such as GTCS [121], TII-SSRC-23 [155], and FLNET2023 [157], and has also been used to transform other existing datasets like ToN_IoT and Bot-IoT [42].
Although a step towards feature standardization is positive in terms of generalizability and transferability, relying a single feature set can also bring downsides. Komisarek et al. [153] criticized CICFlowMeter features as unsuitable for real-time intrusion detection. Sarhan et al. [42] investigated the standardization of NetFlow v9 features for NID and found that they outperform CIC-based features in attack detection.
@@ 5.5.4 Other Observations
From the temporal perspective, 29% of datasets are captured in a single continuous run, 17% are periodic, and 43% are discontinuous – gaps within the data exist – with the latter increasing to 55% since 2020. Similarly to simulation in small networks, we argue that discontinuous captures are easier to perform, allowing flexible adjustments and less strict scheduling. Therefore, they pose a convenient choice for data creators.
@@ 6 Recommendations: Data Selection, Creation, & Usage
After reviewing existing datasets, this section outlines best practices on data-related NID topics: selecting the most appropriate dataset for one’s needs (Section 6.1), creating a custom dataset (Section 6.2), and proper data handling to mitigate biases during experimentation (Section 6.3).
@@ 6.1 Choosing a Dataset
Using a public dataset is the most straightforward and time-efficient approach for NID experimentation. With dozens of such datasets available, the following paragraphs guide the selection of the most suitable ones for specific use cases. However, we refrain from suggesting concrete datasets, as this survey does not aim to assess data quality but rather help others to make informed decisions.
@@
Quality estimation: While we do not assess dataset quality in this paper, we strongly recommend performing it as a part of the data selection process. Despite the previous research on the topic [205, 104, 206], no standardized formal metric for NID data quality has been established within the domain [207]. As “quality” itself is vague, and user needs may vary, formalizing such a general metric is challenging. Therefore, we discuss data quality in a more practical sense and attempt to gauge it by analyzing a) the realism of the data based on the rigorousness of its creation process and b) the contents of the data itself.
A quality dataset should be realistic and error-free. Since documentation (metadata) is one of the few ways to learn about the data creation, its analysis remains a primary resource for assessing realism. Its key details include data purpose, creation process, and related limitations (more in Section 6.2.4). Given this information, datasets with robust traffic generation and realistic environments are preferred over more simplistic ones.
However, our experience shows that analyzing documentation alone is insufficient. Even with decent documentation, numerous datasets were found to contain discrepancies like incorrect entries or timestamps. For this reason, thorough data analysis (e.g., N/A values, incorrect timestamps, packet errors, labeling issues) is essential in addition to documentation analysis to gauge data correctness and realism.
While many studies have highlighted NID data limitations, Flood et al. [8] introduced a practical framework for auditing network data. Given network flows, their approach proposes various metrics to detect mislabeling, artifacts, and poor data diversity. In addition to automated tests, they also conduct several manual checks. Replicating these steps can help identify a dataset’s drawbacks and assess its suitability.
@@ 6.2 Creating a Dataset
In essence, creating a dataset involves ten steps, from defining objectives and designing scenarios to documenting and sharing the data, as shown in Figure 6. In the following paragraphs, we elaborate on the datasets’ desirable properties (Table 4) and recommendations for individual data creation steps to mitigate common dataset limitations outlined in Section 3.2.
Figure 6: The process of creating a NID dataset. The ten steps are further grouped into four phases – preparation, data collection, post-processing, and publication. Ideally, the process is linear, and the data is collected and processed in one run. However, in practice, the process is often iterative, such as re-designing and adding new scenarios or collecting and post-processing the data in multiple turns. Continually-captured datasets have to reiterate steps 2 – 10 periodically by design.
Table 4: Summary of desirable dataset properties. Dataset’s compliance with more properties typically increases data quality. The creation step column outlines which steps during the dataset creation influence the given property. If a dataset needs to stay timely, steps 1 – 10 need to be periodically reiterated.
Property Mentioned in Creation step Description
Correctness [32, 11] 1 – 10 Error-free implementation of the dataset creation steps – e.g., no packet losses, correct feature extraction and labeling, documentation corresponds to the data.
Diversity [33, 14, 105] 2, 3 Diversity and variability of both malicious traffic (multiple launches of the same attack with differing configurations) and benign traffic (rich usage profiles).
Documentation [45, 14, 105] 9 Documentation (metadata) provides detailed and unambiguous information about the dataset.
Extensibility [207] 2, 3, 9, 10 New scenarios, attacks, and traffic capture events can be added on demand.
Label completeness [23, 14, 105, 11] 6 All records are labeled directly, correctly, and unambiguously.
Pre-split [14] 2 – 4 Provides a predefined train-test split to support fair and objective benchmarking.
Privacy preservation [147, 14, 32] 2 – 5, 8 The data does not compromise the privacy (reveal sensitive information) of users. Captures from real networks are properly anonymized.
Realism [17, 144, 32, 54] 1 – 5, 7 The dataset contents and their characteristics reflect real-world networks. The environment is properly and completely configured. The data does not exhibit any unintended properties (e.g., artifacts) as a result of simulation.
Relevance [45, 14, 110] 1 – 6 Dataset scenarios (e.g., network topology, attacks, normal behavior) and the data itself (format, features, labels) are appropriate and relevant for the specified dataset objectives.
Reproducibility [37, 45, 208] 2, 3, 9, 10 The dataset contents can be regenerated on demand.
Timeliness [17, 14, 11] 1 – 3, 10 Contains up-to-date modern attacks and traffic profiles resembling the current threat landscape and the behavior of contemporary computer networks.